OpenStudy (anonymous):

how do you find the y intercept if given the slope = 1/3 and the two coordinate points are (1,5) and (-8,2)how do you find the y intercept if given the slope = 1/3 and the two coordinate points are (1,5) and (-8,2)@Mathematics

OpenStudy (anonymous):

one option is to get the slope-intercept equation(y=mx+b) by plugging in the slope and a coordinate

OpenStudy (anonymous):

so using m=1/3 and the point (1,5) you get 5 = (1/3)(1) + b so solve for b which is b=14/3

OpenStudy (anonymous):

that's right! THNAK YOU!!! MEDAL TO YOU!

OpenStudy (anonymous):

*thank

OpenStudy (anonymous):

so y=1/3x+14/3 in the slope-intercept for b is the y-intercept so (0, 14/3)

OpenStudy (anonymous):

form

OpenStudy (anonymous):

you're welcome!